# Symbolic Music Generation with Non-Differentiable Rule Guided Diffusion

### v2

@@ D.3 Objective Evaluation Setup
Individual Rule Guidance. For each rule, we randomly selected 200 samples from the Muscore test dataset and computed their corresponding rule labels to serve as targets. Subsequently, we generated 200 samples conditioned on each rule label. The rule labels for these generated samples were computed, and the loss between the generated rule label and the target was calculated. The mean and standard deviation of this loss across the 200 samples are presented in Table 3.
@@ G.2 Latent vs Pixel Space
Table 12: Comparison of Average Overlapping Area (OA) for individual rule guidance between diffusion models trained on pixel and latent space.
@@ G.3 Composite Rule Guidance
Weight PH $\downarrow$ ND $\downarrow$ CP $\downarrow$ Quality $\uparrow$
@@
Method $n$ PH $\downarrow$ ND $\downarrow$ CP $\downarrow$ Quality $\uparrow$
@@ H.1 Details
Figure 6: Years of playing music
Figure 7: Years of formal music study

### v4

@@ D.3 Objective Evaluation Setup
Individual Rule Guidance. For each rule, we randomly selected 200 samples from the Muscore test dataset and computed their corresponding rule labels to serve as targets. Subsequently, we generated 200 samples conditioned on each rule label. The rule labels for these generated samples were computed, and the loss between the generated rule label and the target was calculated. For pitch histogram and note density, we use MSE loss. For chord progression, we use 0-1 loss. The mean and standard deviation of this loss across the 200 samples are presented in Table 3.
We also compute OA between the generated samples and the data that comply with the desired rule. Specifically, we cluster 200 generated samples into 4 groups based on note density and compute OA within each group. The mean and standard deviations of OA across the 4 groups are also presented in Table 3.
@@ G.2 Latent vs Pixel Space
Table 13: Comparison of Average Overlapping Area (OA) for individual rule guidance between diffusion models trained on pixel and latent space. OA is computed using the full dataset as reference.
@@ G.3 Composite Rule Guidance
Weight PH $\downarrow$ ND $\downarrow$ CP $\downarrow$ OA full $\uparrow$
@@
Method $n$ PH $\downarrow$ ND $\downarrow$ CP $\downarrow$ OA full $\uparrow$